how can i find the answer? Dave has eight more dimes than nickels in his pocket. In total, the coins in his pocket are
worth $3.05. How many of each type of coin does he have?

Answer by ikleyn(52781)

.
how can i find the answer? Dave has eight more dimes than nickels in his pocket.
In total, the coins in his pocket are worth $3.05. How many of each type of coin does he have?
~~~~~~~~~~~~~

Let x be the number of nickels; then the number of dimes is (x+8).


The value of nickels is 5x cents; the value of dimes is 10*(x+8) cents.


The total value is $3.05, or 305 cents; so we write this equation

    5x + 10*(x+8) = 305.


Simplify and find x

    5x + 10x + 80 = 305

       15x = 305 - 80

       15x = 225

         x = 225/15 = 15.


ANSWER.  15 nickels and 15+8 = 23 dimes.


CHECK.  The total value is 15*5 + 23*10 = 75 + 230 = 305 cents = $3.05.   ! correct !

Solved.